# Dark Sector Tunneling Field Potentials for a Dark Big Bang

### v1

@@ VI Conclusion
We have shown that nearly all of the important characteristics describing a DBB can be put in terms of the tunneling field parameters $m$, $\mu$, and $\lambda$. Even the gravity waves released from the PT, one of the most important observables of the DBB, can be determined from the tunneling field parameters (Eq. 31). Evidence for the existence of a stochastic background of gravity waves permeating the universe was first detected by NanoGrav in June 2023 [12]. As research on gravity waves continues to advance, signatures of a DBB may become differentiable from the stochastic background. In fact, signatures of the Dark Big Bang may have already been observed in NanoGrav’s 15 year data set [1].
In the future, it would be interesting to see how the projected sensitivities of upcoming gravity wave surveys translate onto the parameter space of the tunneling field potential. Future research could also consider weak couplings between the dark and visible sectors through portals. These couplings could give rise to additional observable signatures of the DBB [1] and new bounds on the tunneling field potential parameters.
@@ Acknowledgments
RC would like to thank Saiyang Zhang for useful discussions during the later stages of this project.
@@ Appendix A Analytic Behavior
In this section we discuss the analytic behavior of the phase transition parameters. We will separate the parameter space into two regimes: Region 1 ($\mu\gtrsim$ m) and Region 2 ($\mu\gg$ m).
@@ A.1 Region 1
S has a singularity at $\mu^{2}=2\lambda m^{2}$, which is the driving factor for the behavior of $\alpha$ and $T_{*}$ for Dark Big Bangs where $m$ and $\mu$ are comparable in size. Since $\mu>\sqrt{2\lambda}m$, we are right above the singularity in this region of parameter space. Setting $\lambda=1$, $\alpha$ scales like:
@@ A.2 Region 2
Another allowed parameter space for a DBB consistent with the imposed bounds opens up when $\mu\gg m$ for reasons that will be shown here. In this region we are far from the singularity. It is convenient to rewrite the tunneling potential and PT parameters in terms of the variable $\frac{m^{2}}{\mu^{2}}$. In the following analysis, we set higher order terms of $\frac{m^{2}}{\mu^{2}}$ to zero and set $\lambda=1$. To first-order approximation, the Euclidean action scales like
@@
$\Delta V\sim\mu^{4}(1-\frac{8}{3}\frac{m^{2}}{\mu^{2}})$ (36)
@@
These scalings show that region 2 of the parameter space should be expected. Near the singularity, $\alpha$ decreases from positive infinity, shooting quickly through the allowed parameter space. Then, if $m$ is fixed and $\mu$ increased, $\alpha$ will return to allowed values from below the lower bound.
We can also find the slope of the bounds on the DBB when $\mu\gg m$. Writing the lower bound on $\alpha$ (CMB $\Delta N_{\rm{eff}}$) as a function of $\mu$, we find
@@
For the upper bound on $\alpha$ (Dark Matter), we find
@@
These scaling relations are reflected in the parameter space where the upper bound on $\alpha$ has a steeper slope than the lower bound (see Fig. 5 or Fig. 7). The steepest slope is the temperature (Lyman-$\alpha$) bound which has the scaling $\mu\sim m^{2}$, as can be seen from Eq. 38.
@@ Appendix B Critical Point Evaluation
The purpose of this section is to give numerical values for the tunneling field potentials and phase transition parameters for critical points tabulated in Tab. 1. Since the critical points found are numerical approximations for the boundaries of the parameter space, some values calculated violate the bounds imposed on the DBB and some fall just within the parameter space. Tab. 3 is meant to give a general idea of the tunneling field and PT parameters near the critical points. For all calculations, we set $\lambda=1$.
R2 Opens R1 Opens Intersection PS Closes (TCC) PS Closes
@@
$\Delta\phi$ [MeV] $1.58\times 10^{-4}$ $2.03\times 10^{-4}$ $1.15\times 10^{10}$ $1.60\times 10^{12}$ $1.53\times 10^{30}$
@@
Table 3: Tunneling potential and phase transition parameters for the critical points in the $m-\mu$ parameter space. To calculate $T_{\rm{DS},*}$, we fixed $g_{\rm{DS}}=2$. We have also included the maximum mass of dark matter candidates produced in each DBB scenario, according to Eq. 25. For the critical point where Region 2 (R2) opens (column 2), T is slightly below the Lyman-$\alpha$ bound. Notably, for both critical points where the parameter space (PS) closes (columns 5 & 6), $T_{\rm{DS},*}>T_{*}$ yet $\alpha$ falls within the bounds (it is typical for the dark sector reheating temperature to be less than the visible sector temperature at the DBB). This is a result of our choice for $g_{\rm{DS}}$, which is much smaller than $g_{\rm{eff}}$ at high temperatures.

### v3

@@ VI Conclusion and Future Directions
We have shown that nearly all of the important characteristics describing a DBB can be put in terms of the tunneling field parameters $m$, $\mu$, and $\lambda$. Even features of the gravity waves released from the PT, one of the most important observables of the DBB, can be determined from the tunneling field parameters (Eq. 32). Evidence for the existence of a stochastic background of gravity waves permeating the universe was first detected by NanoGrav in June 2023 [12]. As research on gravity waves continues to advance, signatures of a DBB may become differentiable from the stochastic background. In fact, signatures of the Dark Big Bang may have already been observed in NanoGrav’s 15 year data set [1]. In the case of Dark WIMPs, evidence for a Dark Big Bang can be observed through a simultaneous GW and CMB signal [1]. The dark radiation density present in the dark sector contributes to $\Delta N_{\rm{eff}}$, which can be measured in the upcoming CMB observations like CMB-S4 [11] and the Simons Observatory [10].
In the future, it would be interesting to see how the projected sensitivities of upcoming gravity wave surveys translate onto the parameter space of the tunneling field potentials. Future research could also consider weak couplings between the dark and visible sectors through portals. These couplings could give rise to additional observable signatures of the DBB [1] and new bounds on the tunneling field potential parameters.
@@ Acknowledgments
R.C. would like to thank Saiyang Zhang for useful discussions during the later stages of this project. C.I. acknowledges funding from Colgate University via the Research Council and the Picker Interdisciplinary Science Institute.
@@ Appendix A Analytic Behavior
In this section we discuss the analytic behavior of the phase transition parameters to clarify why the parameter space is expected to look as it does. We will separate the parameter space into two regimes: Region 1 ($\mu\gtrsim$ m) and Region 2 ($\mu\gg$ m).
@@ A.1 Region 1
S has a singularity at $\mu^{2}=2\lambda m^{2}$, which is the driving factor for the behavior of $\alpha$ and $T_{*}$ for Dark Big Bangs where $m$ and $\mu$ are comparable in size. When $\mu\gtrsim\sqrt{2\lambda}m$, we are right above the singularity in this region of parameter space. Setting $\lambda=1$, $\alpha$ scales like:
@@ A.2 Region 2
Another allowed parameter space for a DBB consistent with the imposed bounds opens up when $\mu\gg m$ for reasons that will be analytically motivated here. When $\mu\gg m$, we are far from the singularity in $S$. It is convenient to rewrite the PT parameters in terms of the variable $\frac{m^{2}}{\mu^{2}}$. In the following analysis, we set higher order terms of $\frac{m^{2}}{\mu^{2}}$ to zero and set $\lambda=1$. To first-order approximation, the Euclidean action scales like
@@
$\Delta V\sim\mu^{4}(1-\frac{4}{3}\frac{m^{2}}{\mu^{2}})$ (37)
@@
These scalings show that region 2 of the parameter space should be expected. Near the singularity (Region 1), $\alpha$ decreases from positive infinity, shooting quickly through the allowed parameter space. Then, if $m$ is fixed and $\mu$ increased, $\alpha$ will return to allowed values from below the lower bound.
We can also find the slope, in $m$-$\mu$ space, of the bounds on the DBB when $\mu\gg m$. Writing the lower bound on $\alpha$ (Dark Matter) as a function of $\mu$, we find
@@
For the upper bound on $\alpha$ (CMB $\Delta N_{\rm{eff}}$), we find
@@
These scaling relations are reflected in the parameter space where the border of the upper bound (blue region) on $\alpha$ has a steeper slope than the border of the lower bound (orange region; see Fig. 5 or Fig. 8). The steepest slope is the temperature (Lyman-$\alpha$) bound (green region) which has the scaling $\mu\sim m^{2}$, as can be seen from Eq. 39.
@@ Appendix B Critical Point Evaluation
The purpose of this section is to give numerical values for the tunneling field potentials and phase transition parameters for the critical points tabulated in Tab. 1. Since the critical points found are numerical approximations for the boundaries of the parameter space, some values calculated violate the bounds imposed on the DBB and some fall just within the parameter space. Tab. 3 is meant to give a general idea of the tunneling field and PT parameters near the critical points. For all calculations, we set $\lambda=1$.
R2 Opens R1 Opens Regions Merge PS Closes (TCC) PS Closes
@@
$\Delta\phi$ [MeV] $1.58\times 10^{-4}$ $2.03\times 10^{-4}$ $1.15\times 10^{10}$ $1.60\times 10^{12}$ $1.53\times 10^{20}$
@@
Table 3: Tunneling potential and phase transition parameters for the critical points in the $m$-$\mu$ parameter space. To calculate $T_{\rm{DS},*}$, we fixed $g_{\rm{DS}}=2$. We have also included the maximum mass of dark matter candidates produced in each DBB scenario, according to Eq. 26. For the critical point where Region 2 (R2) opens (column 2), $T_{*}$ is slightly below the Lyman-$\alpha$ bound. Notably, for both critical points where the parameter space (PS) closes (columns 5 & 6), $T_{\rm{DS},*}>T_{*}$ yet $\alpha$ falls within the bounds (it is typical for the dark sector reheating temperature to be less than the visible sector temperature at the DBB). This is a result of our choice for $g_{\rm{DS}}$, which is much smaller than $g_{\rm{eff}}$ at high temperatures.